A leaf hangs from a branch 12 feet in the air. It falls to the ground at a rate of 0.25 feet per second. Which graph could represent the leaf’s height in feet as a function of time, in seconds, after leaving the branch?


1
Expert's answer
2020-10-21T17:15:15-0400

The equation of a straight line is given as "y=mx+c"

In this case, y is the leaf's height in feet, x is the time in seconds, c is the initial height of 12 feet and m is the rate at which the leaf falls.

Since the leaf is falling, the height is decreasing at the rate of 0.25 feet per second

In this case:

"y=" leaf's height in feet

"m=-0.25"

"x=" time in seconds

"c=12"

Therefore the graph is that of the straight line whose equation is

"y=12+(-0.25)*x"

"y=12-0.25x"